### v1

# LW-FedSSL: Resource-efficient Layer-wise Federated Self-supervised Learning

Abstract
Many recent studies integrate federated learning (FL) with self-supervised learning (SSL) to take advantage of raw training data distributed across edge devices. However, edge devices often struggle with high computation and communication costs imposed by SSL and FL algorithms. To tackle this hindrance, we propose LW-FedSSL, a layer-wise federated self-supervised learning approach that allows edge devices to incrementally train one layer of the model at a time. LW-FedSSL comprises server-side calibration and representation alignment mechanisms to maintain comparable performance with end-to-end FedSSL while significantly lowering clients’ resource requirements. The server-side calibration mechanism takes advantage of the resource-rich server in an FL environment to assist in global model training. Meanwhile, the representation alignment mechanism encourages closeness between representations of FL local models and those of the global model. Our experiments show that LW-FedSSL has a $\times3$ lower memory requirement and a 3.2×3.2$\times$3.2 × cheaper communication cost than its end-to-end counterpart. We also explore a progressive training strategy called Prog-FedSSL that outperforms end-to-end training with a similar memory requirement and a 1.8×1.8$\times$1.8 × cheaper communication cost.
1 Introduction
A significant portion of real-world data, valuable for practical AI applications, resides on edge devices. Such data is often unlabeled, making the adoption of self-supervised learning (SSL) strategies Chen et al. (2020); He et al. (2020); Chen and He (2021) essential. SSL enables models to learn from raw data by generating their own supervisory signals, eliminating the need for labels. Most SSL methods follow centralized training schemes, often requiring an exhaustive data collection process to build a centrally stored dataset. Although some data, such as natural images, can be conveniently collected, it can raise significant concerns for privacy-sensitive data, such as medical data. Meanwhile, decentralized learning approaches like federated learning (FL) Konečnỳ et al. (2016); McMahan et al. (2017) enables privacy-preserving collaborative model training. An FL system functions by collecting trained model parameters from edge devices (a.k.a., clients) rather than data. This mechanism safeguards the confidentiality of private data held by different parties. As a result, federated self-supervised learning (FedSSL) approaches have recently emerged as a potential strategy for leveraging raw data in distributed environments Zhuang et al. (2021); Li et al. (2023); Wang et al. (2023).
SSL, especially when involving state-of-the-art models like transformers Vaswani et al. (2017); Dosovitskiy et al. (2021a), can impose substantial computation demands. Unfortunately, clients in an FL environment often operate with limited computation and communication resources. Therefore, it is crucial to explore resource-efficient SSL approaches for FL systems. A straightforward way to reduce computation demands on low-memory devices is to conduct the training with a smaller batch size. However, using a small batch size for SSL approaches can diminish the quality of learned representations and compromise performance Chen et al. (2020); Li et al. (2023). Moreover, simply reducing the batch size does not lower the FL communication cost since clients still need to exchange the full model with the server. On the other hand, using a large batch size would exclude low-memory devices from participating in the FL process.
One promising solution is the layer-wise training approach, where a model is systematically trained one layer at a time, gradually increasing its depth. The term “layer” can refer to either an individual layer or a block comprising multiple layers. In layer-wise training, only a single layer is active at a time for training, significantly reducing computation requirements. Moreover, only the active layer needs to be exchanged with the server, further reducing communication costs. Table 1 shows a significant reduction in a client’s resource requirements for layer-wise training compared to end-to-end training.
Memory
Comm.
Accuracy
FedMoCoV3 15601 7352.49 54.99
LW-FedMoCoV3 3927 613.19 46.41
Table 1: FedMoCoV3 (end-to-end) vs. LW-FedMoCoV3 (layer-wise). We use a ViT-Tiny backbone with MoCo v3 as the SSL approach. The FL process involves 10 clients using a batch size of 1024 and the number of communication rounds set to 180. The unlabeled portion of STL-10 dataset is used to setup the clients’ data, while the labeled training and testing portions are used for linear evaluation. We compare resource requirements for a single client.
Although layer-wise training significantly reduces resource requirements, it may not be able to match the performance of end-to-end training as shown in Table 1. In each stage of layer-wise training, a new layer is sequentially attached to the model, and all preceding layers are frozen, excluding them from further updates. Given a fixed number of FL rounds for each stage, it can be challenging to train the newly added layer to work seamlessly with the previous layers, especially if the number of rounds is small. Although increasing the number of rounds could potentially alleviate this problem, it would impose extra communication and computation overhead on the clients, canceling the benefits of layer-wise training.
To address these challenges, we propose LW-FedSSL, by introducing two simple, yet effective mechanisms: server-side calibration and representation alignment. Although clients in an FL environment may have limited resources, the central server typically possesses significant computation power. Therefore, in LW-FedSSL, the server-side calibration mechanism leverages the computation resources of the central server to train the global model on an auxiliary dataset. This mechanism solely operates on the central server and incurs no additional overhead on the clients. Additionally, the representation alignment mechanism encourages alignment between local models and the server-side-trained global model during the local training process.
Our contributions can be summarized as follows: (i) We conduct an in-depth exploration of layer-wise training for FedSSL. (ii) We introduce LW-FedSSL, a framework designed to improve the effectiveness of layer-wise training in the context of FedSSL. LW-FedSSL achieves comparable performance to end-to-end training while significantly lowering the resource requirements of FL clients. (iii) We also explore a progressive training approach called Prog-FedSSL. Although the progressive approach requires more resources than the layer-wise approach, it still remains more resource-efficient than end-to-end training while significantly outperforming both.
2 Preliminary
Federated Learning.
The goal of FL is to train a global model $M$, by learning from the data residing on clients while ensuring their privacy. This objective can be expressed as
$M^{*}=\operatorname*{arg\,min}_{M}\sum_{i=1}^{N}w^{i}\mathcal{L}(M,D^{i}),$ (1)
where $N$ is the number of clients and $\mathcal{L}$ is the local loss. Here, $w^{i}=|D^{i}|/|D|$, which is the weight term of $i$-th client, with $|D^{i}|$ denoting the number of samples in local dataset $D^{i}$ and $|D|$ denoting the total number of samples in $D=\bigcup_{i=1}^{N}D^{i}$. An FL communication round consists of the following four key steps. At the beginning of each round, (i) the server distributes a base global model $M$ to all participating clients. (ii) Each client $i\in[1,N]$ performs training on its local dataset $D^{i}$ to produce a local model $M^{i}$. After the training, (iii) local models are transmitted back to the server. Finally, (iv) the sever aggregates the received local models through weighted averaging McMahan et al. (2017) to update the global model: $M\leftarrow\sum^{N}_{i=1}w^{i}M^{i}$. These steps are repeated for a number of rounds $R$. In the context of SSL, the local dataset $D^{i}$ is presumed to be unlabeled, and thus, clients execute SSL to train a model $M^{i}$ consisting of an encoder $F^{i}$ and its MLP heads, $H^{i}$ and $P^{i}$.
Self-supervised Learning.
We select MoCo v3 Chen et al. (2021) as the SSL approach in our work. MoCo v3 features a Siamese network structure with an actively trained online branch, and a target branch, which is a moving-averaged copy of the online branch. The online branch consists of an encoder $F$, a projection head $H$, and a prediction head $P$. Meanwhile, the target branch consists of a momentum encoder $F_{k}$ and a momentum projection head $H_{k}$. Given an input sample $x$, it undergoes augmentation, creating two views, $x_{q}$ and $x_{k^{+}}$. These views are then fed into the online and target branches, respectively, to obtain a vector $q$ and its corresponding positive pair, $k^{+}$. Likewise, negative pairs, $k^{-}$, are derived from other samples within the same batch through the target branch. The online branch is trained using the InfoNCE loss Van den Oord et al. (2018),
$\ell(q,k,\tau)=-\log\frac{\exp(q\cdot k^{+}/\tau)}{\exp(q\cdot k^{+}/\tau)+\sum_{j=1}^{K}\exp(q\cdot k_{j}^{-}/\tau)},$ (2)
where $k=(k^{+},\{k_{j}^{-}\}_{j=1}^{K})$ and $\tau$ is the temperature parameter. Using a momentum parameter $\mu$, the target branch is updated as: $F_{k}\leftarrow\mu F_{k}+(1-\mu)F$ and $H_{k}\leftarrow\mu H_{k}+(1-\mu)H$.
Figure 1: Local training process of LW-FedMoCoV3 across different stages for $i$-th client. Each stage increases the depth of the encoder by adding a new layer. During stage $s$, only the corresponding layer $L_{s}$ is actively updated.
Figure 2: LW-FedSSL at stage $s$. The server-side calibration mechanism trains all existing layers in the current sub-model. The representation alignment mechanism aligns the representations of the local model with those of the global model.
3 Related Work
Layer-wise training was initially introduced in the context of deep belief networks and restricted Boltzmann machines Hinton et al. (2006); Bengio et al. (2006). The original motivation behind its development was to address the challenges associated with training deep neural networks, such as vanishing gradients. While layer-wise training has become less common over the years, it can still be effectively employed as a resource-efficient strategy within FL systems. Training the model incrementally, one layer at a time, offers a distinct resource-saving advantage over end-to-end training. To date, only a limited number of studies Wang et al. (2022); Guo et al. (2023) have explored layer-wise training in the FL context, and even fewer Guo et al. (2023) have examined its application in the domain of FedSSL.
Some studies Caldas et al. (2018); Horvath et al. (2021) reduce clients’ computation costs by selectively dropping parts of the model. Federated Dropout Caldas et al. (2018) randomly removes a subset of neurons or filters from a layer to reduce the computation cost while relying on compression techniques to minimize the communication overhead. FjORD Horvath et al. (2021), a similar approach, introduces an ordered dropout mechanism instead of random dropping. Meanwhile, Wang et al. Wang et al. (2022) propose a progressive training approach for FL called ProgFed. While similar to layer-wise training in gradually increasing the model depth, ProgFed differs by training all existing layers at each stage. The aforementioned studies primarily focus on federated supervised learning tasks that require labeled data for training.
On the other hand, Huo et al. Huo et al. (2021) study layer-wise SSL in the speech domain. While the study was intended for an on-device training scenario, the authors mainly examined it in a centralized setting. Another study by Guo et al. Guo et al. (2023) is closely related to our work, as it also explores layer-wise training for FedSSL. The authors propose a depth dropout technique which aims to reduce both computation and communication overhead but may lead to a slight decrease in performance. In contrast to the above studies, our work offers an in-depth exploration of layer-wise training in FedSSL. However, as we have previously demonstrated, layer-wise training alone may not achieve the same level of performance as end-to-end training. To bridge this gap, we introduce LW-FedSSL, a layer-wise training approach that can compete end-to-end training with significantly lower client resource requirements. Moreover, we explore a progressive training approach called Prog-FedSSL, which outperforms end-to-end training while requiring fewer resources.
4 Proposed Methods
4.1 LW-FedMoCoV3
Firstly, we discuss the pure layer-wise training baseline, referred to as LW-FedMoCoV3. Figure 1 illustrates the local training process of LW-FedMoCoV3 across different stages. For an encoder $F$ with a total of $S$ layers, the training process is divided into $S$ stages. Each stage $s\in[1,S]$ runs for a fixed number of FL communication rounds, $R_{s}$. The training process starts with an empty encoder $F$, and each stage $s$ sequentially adds a new layer $L_{s}$ to $F$, gradually increasing its depth. During stage $s$, only the corresponding layer $L_{s}$ is updated, while all prior layers $L_{1}$ to $L_{s-1}$ within $F$ are frozen, serving only for inference. This strategy saves memory and computation, benefiting resource-constrained devices in FL environments. Moreover, in each communication round, clients are only required to send the active layer $L_{s}$ (and the MLP heads) to the server for aggregation. Likewise, the server only needs to broadcast the aggregated layer $L_{s}$ back to clients. This reduces both upload and download communication costs for clients. Although layer-wise training offers better resource efficiency than end-to-end training, there remains a notable performance gap in linear evaluation, as we have shown in Table 1.
4.2 LW-FedSSL
To address the performance gap between layer-wise and end-to-end training for FedSSL, we propose LW-FedSSL by introducing two crucial mechanisms: server-side calibration, integrated at the server, and representation alignment, integrated into the local training process. Figure 2 illustrates these two mechanisms of LW-FedSSL at stage $s$.
Server-side calibration: In layer-wise training, only the most recently added layer, $L_{s}$, is actively trained by clients at any given stage $s$. This may limit prior layers’ ability to adapt to the newly added layer, potentially degrading the model’s performance. Given a limited number of rounds $R_{s}$ available for each stage, solely training $L_{s}$ may prove insufficient to match the training received by the end-to-end strategy. On the other hand, increasing the number of rounds could introduce additional resource demands on clients. To tackle this, we can take advantage of resources available at the FL server. The server can contribute by training the current sub-model on a small auxiliary dataset, $D^{g}$, at the end of every communication round. Specifically, following the aggregation process, the server can train the aggregated encoder $F$ across all existing layers (i.e., $L_{1}$ to $L_{s}$) before broadcasting $F$ back to the clients. The dataset $D^{g}$ may originate from a different domain than that of the clients’ data, as the server cannot directly access data within clients. Furthermore, $D^{g}$ may typically be unlabeled, requiring server-side training to also utilize an SSL approach. We believe that training across all layers in the current sub-model plays a crucial role in calibrating the layers to function harmoniously. Also, our assumption regarding the auxiliary data at the server can be reasonably met since the server may require a small dataset for various purposes, such as performance monitoring, pre-training, distillation, or other tasks in the FL context Li et al. (2023); Sattler et al. (2021).
Algorithm 1 LW-FedSSL (Server-side)
Input: encoder $F$, projection head $H$, prediction head $P$, auxiliary dataset $D^{g}$, number of epochs $E$, momentum $\mu$, temperature $\tau$, number of stages $S$, number of rounds per stage $R_{s}$, ${R\leftarrow\sum_{s=1}^{S}R_{s}}$
Output: encoder $F$
1:Server executes:
2:Initialize stage: $s=0$
3:for round $r=1,2,\dots,R$ do
4:     if $r>\sum_{j=1}^{s}R_{j}$ then
5:         $s\leftarrow s+1$
6:         Append a new layer $L_{s}$ to $F$
7:     Unfreeze $F$
8:     $F,H,P\leftarrow\text{Server\_SSL}(D^{g},E,\mu,\tau,F,H,P)$
9:     Freeze $F$ up to layer $L_{s-1}$
10:     for client $i=1,2,\dots,N$ in parallel do
11:         $L_{s}^{i},H^{i},P^{i}\leftarrow\text{Local\_SSL}(i,F,H,P)$ Algo. 2
12:         $w^{i}=\frac{|D^{i}|}{|D|}$, where $D=\bigcup_{i=1}^{N}D^{i}$
13:     $L_{s},H,P\leftarrow\text{Aggregation}(\{L^{i}_{s},H^{i},P^{i},w^{i}\}^{N}_{i=1})$
14:     Update $L_{s}$ in $F$
15:return $F$
Algorithm 2 LW-FedSSL (Client-side)
Input: encoder $F$, projection head $H$, prediction head $P$, local dataset $D^{i}$, number of local epochs $E$, momentum $\mu$, temperature $\tau$, weight term $\alpha$
Output: layer $L^{i}_{s}$, local projection head $H^{i}$, local prediction head $P^{i}$
1:Client executes:
2:Local_SSL($i,F,H,P$):
3:Online branch: $F^{i}\leftarrow F$ and $H^{i}\leftarrow H$ and $P^{i}\leftarrow P$
4:Target branch: $F^{i}_{k}\leftarrow F$ and $H^{i}_{k}\leftarrow H$
5:for epoch $e=1,2,\dots,E$ do
6:     for each batch $x\in D^{i}$ do
7:         $x_{1}\leftarrow\text{{Augment}}(x)$ and $x_{2}\leftarrow\text{{Augment}}(x)$
8:         $z_{1}\leftarrow F(x_{1})$ and $z_{2}\leftarrow F(x_{2})$
9:         $z_{1}^{i}\leftarrow F^{i}(x_{1})$ and $z_{2}^{i}\leftarrow F^{i}(x_{2})$
10:         $q_{1}^{i}\leftarrow P^{i}(H^{i}(z_{1}^{i}))$ and $q_{2}^{i}\leftarrow P^{i}(H^{i}(z_{2}^{i}))$
11:         $k_{1}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{1}))$ and $k_{2}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{2}))$
12:         $\ell_{\text{con}}\leftarrow\ell(q^{i}_{1},k^{i}_{2},\tau)+\ell(q^{i}_{2},k^{i}_{1},\tau)$ Eq. (2)
13:         $\ell_{\text{align}}\leftarrow\ell(z^{i}_{1},z_{2},\tau)+\ell(z^{i}_{2},z_{1},\tau)$ Eq. (3)
14:         $\mathcal{L}\leftarrow\ell_{\text{con}}+\alpha\ell_{\text{align}}$
15:         $F^{i},H^{i},P^{i}\leftarrow\text{Update}(\nabla\mathcal{L})$
16:         $F^{i}_{k},H^{i}_{k}\leftarrow\text{Momentum\_Update}(F^{i},H^{i},\mu)$
17:$L_{s}^{i}\leftarrow\text{Get active layer from }F^{i}$
18:return $L_{s}^{i},H^{i},P^{i}$
Representation alignment: FL clients often exhibit distinct local data distributions resulting from diverse device usage patterns and application settings. Such data heterogeneity can introduce bias to the local training process, negatively affecting the performance of the global model. To mitigate this, some studies Li et al. (2021) encourage local training processes across different clients to converge toward parameters that are closely aligned with the global parameters. Often, this is achieved by aligning the representations of local models with those of the global model while simultaneously pushing them away from old representations (from old local models of previous FL rounds). In contrast, our representation alignment mechanism does not require old local parameters. As a result, clients no longer need to store and perform inference on old parameters, reducing the overhead. Given a positive pair of views, $x_{q}$ and $x_{k^{+}}$, our approach uses local encoder $F^{i}$ and the global encoder $F$ to extract local representation $z^{i}=F^{i}(x_{q})$ and positive global representation $z^{+}=F(x_{k^{+}})$. Negative global representations $z^{-}$ are extracted from other samples. Similar to Equation 2, we can calculate the alignment loss as
$\ell(z^{i},z,\tau)=-\log\frac{\exp(z^{i}\cdot z^{+}/\tau)}{\exp(z^{i}\cdot z^{+}/\tau)+\sum_{j=1}^{K}\exp(z^{i}\cdot z_{j}^{-}/\tau)},$ (3)
where $z=(z^{+},\{z_{j}^{-}\}^{K}_{j=1})$. This loss can be integrated into local training as an auxiliary loss for regularization, with a weight term $\alpha$ to control its contribution. The representation alignment mechanism also helps local models stay close to the global model that has gone through training at the server.
Integrating the aforementioned mechanisms improves the effectiveness of layer-wise training, enabling LW-FedSSL to perform comparably with end-to-end training, all the while substantially reducing resource demands on FL clients. The details of LW-FedSSL are presented in Algorithms 1 and 2. Algorithm 1 outlines the server-side execution, which handles gradual increase in the model depth based on stage, server-side calibration mechanism, and aggregation. Each client executes Algorithm 2, performing local SSL training with representation alignment for regularization.
Cost Linear Evaluation Fine-tune
Memory FLOPs Comm. CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
MoCo v3 70.82 $\pm$ 0.06 46.09 $\pm$ 0.45 63.84 $\pm$ 0.42 28.61 $\pm$ 0.45 86.27 $\pm$ 0.26 64.57 $\pm$ 0.38 69.78 $\pm$ 0.91 43.95 $\pm$ 0.48
Scratch 23.47 $\pm$ 1.20 5.83 $\pm$ 0.42 17.50 $\pm$ 1.00 2.68 $\pm$ 1.09 63.23 $\pm$ 0.94 41.69 $\pm$ 0.58 44.08 $\pm$ 0.13 27.99 $\pm$ 0.33
FedMoCoV3 1.00×1.$0$ 1.00×1.$0$ 1.00×1.$0$ 60.64 $\pm$ 0.14 34.66 $\pm$ 0.42 54.96 $\pm$ 0.03 20.55 $\pm$ 0.28 79.46 $\pm$ 0.32 58.70 $\pm$ 0.32 62.59 $\pm$ 0.36 40.62 $\pm$ 0.28
FedBYOL 1.00×1.$0$ 1.00×1.$0$ 1.00×1.$0$ 55.38 $\pm$ 0.03 28.64 $\pm$ 0.04 51.01 $\pm$ 0.07 16.34 $\pm$ 0.05 77.25 $\pm$ 0.11 56.92 $\pm$ 0.39 60.36 $\pm$ 0.18 39.58 $\pm$ 0.39
[…]
Table 2: Linear evaluation and fine-tuning performance of different approaches. We conduct the experiments three times with different seeds. Three best values, except for the centralized MoCo v3, are marked in bold. We compare the memory, FLOPs, and communication costs (download + upload) for a local training process. For FedSimCLR, we use a batch size of 512.
(a) Memory.
(b) FLOPs.
(c) Download.
(d) Upload.
Figure 3: Computation and communication resources required for a local training process.
4.3 Prog-FedSSL
Inspired by Wang et al. Wang et al. (2022), we explore progressive training for FedSSL, denoted as Prog-FedSSL. Similar to LW-FedMoCoV3, progressive training also starts with an empty encoder $F$, appending a new layer $L_{s}$ to $F$ at each stage $s$. However, in contrast to updating only a single layer $L_{s}$, progressive training updates all existing layers (i.e., $L_{1}$ to $L_{s}$) during each stage $s$. This distinction results in higher computation and communication costs for Prog-FedSSL than layer-wise training approaches. Nonetheless, when compared with end-to-end training, which requires clients to train a full model with $S$ layers (i.e., $L_{1}$ to $L_{S}$) at each round, Prog-FedSSL enables clients to train only a sub-model with $s$ layers $(s\leq S)$. Moreover, clients only need to exchange the sub-model with the server. In essence, Prog-FedSSL saves both computation and communication costs for $(S-s)$ layers at each stage $s$ compared to end-to-end training. As we will demonstrate in Section 5.2, progressive training outperforms both layer-wise and end-to-end training approaches. Nevertheless, it is important to note that the memory requirement of progressive training gradually increases as the stages progress, becoming equivalent to that of end-to-end training at the final stage (i.e., when $s=S$). This may limit low-memory devices from participating in the FL process, whereas the memory requirement of layer-wise training is significantly lower due to training a single layer $L_{s}$ at a time.
5 Experiment
5.1 Experimental setup
[…]
For performance comparison, we primarily use linear evaluation Kolesnikov et al. (2019); Bachman et al. (2019); Chen et al. (2020) in which a linear classifier is trained on top of the frozen encoder $F$. The classifier is trained for 20 epochs using a batch size of 256. We use the AdamW optimizer with a base learning rate of 3e-2 and a weight decay of 1e-5. The learning rate decays with a cosine schedule. We use the RandAugment function available in the PyTorch Paszke et al. (2019) framework for augmentation. For fine-tuning, we set the base learning rate to 1e-3 and the number of epochs to 40 with a warmup period of 10 epochs. Other settings are similar to those of linear evaluation.
Figure 4: Linear evaluation performance of LW-FedSSL, the baseline LW-FedMoCoV3, and its two variants incorporating either server-side calibration or representation alignment. Weight transfer between blocks is not performed for these two variants, as they demonstrate better performance without it.
5.2 Performance and Resource Comparison
Table 2 compares the performance of different approaches. In terms of linear evaluation, LW-FedSSL demonstrates comparable performance to the end-to-end training of FedMoCoV3 and significantly outperforms the purely layer-wise training of LW-FedMoCoV3. LW-FedSSL also outperforms or holds a competitive position against FedMoCoV3 in fine-tuning. The progressive training approach, Prog-FedSSL, outperforms all other approaches in both linear evaluation and fine-tuning performance. Interestingly, the fine-tuning performance of FedMoCoV3 is even lower than that of LW-FedMoCoV3 in some cases. Please note that we apply the same fine-tuning settings, which may not be optimal for all approaches.
In Figure 3, we present resource requirements for a single client. Figure 3a measures the memory usage at each local epoch, showing that the memory requirement of LW-FedSSL is $\times3$ lower than that of FedMoCoV3 at peak usage. The memory usage of Prog-FedSSL gradually increases and eventually matches that of FedMoCoV3 in the final stage. Figure 3b shows that LW-FedSSL also maintains a lower FLOPs requirement than FedMoCoV3, with 2.1×2.1$\times$2.1 × lower total FLOPs consumption. Additional memory and FLOPs requirements of LW-FedSSL over LW-FedMoCoV3 arise from the representation alignment mechanism, which requires inference on the global model to obtain global representations.
In Figures 3c and 3d, we investigate the communication cost required for model exchange (download and upload) between the server and a client. Note that we only consider the encoder part for comparison, excluding the projection and prediction heads, which are common to all approaches with constant communication costs. In FedMoCoV3, a client needs to exchange the full encoder with the server at each communication round. Meanwhile, in LW-FedMoCoV3, a client only needs to exchange the active layer (i.e., $L_{s}$). Thus, both FedMoCoV3 and LW-FedMoCoV3 result in constant download and upload costs across different communication rounds. For Prog-FedSSL, both download and upload costs gradually increase as the number of active layers (i.e., $L_{1}$ to $L_{s}$) grows with the stage $s$. As for LW-FedSSL, a client needs to download layers $L_{1}$ to $L_{s}$ from the server, as these are updated through the server-side calibration mechanism. However, the client uploads a single layer $L_{s}$, as only it is updated by the local training process. Therefore, the download cost of LW-FedSSL increases with each step $s$, while the upload cost remains constant. Upon comparing the total communication cost over 180 rounds, LW-FedSSL is 1.8×1.8$\times$1.8 × cheaper in terms of download cost and 12×12$\times$12 × cheaper in terms of upload cost compared to FedMoCoV3.
5.3 Ablation Study
Figure 5: Impact of batch size on different approaches. (a) Linear evaluation results on the STL-10 dataset with different batch sizes. (b) Peak memory requirements for different batch sizes.
In Figure 4, we compare LW-FedSSL with the baseline LW-FedMoCoV3 and its two variants, which solely incorporate either server-side calibration or representation alignment. Incorporating server-side calibration improves performance on the CIFAR-10 dataset (from which the auxiliary data is sampled), but leads to a decline on other datasets. This may result from training on auxiliary data, whose update direction conflicts with that of local training. Compared to the baseline, incorporating representation alignment obtains better performance on the STL-10 dataset, with comparable performance on the others. Meanwhile, LW-FedSSL, which incorporates both mechanisms, consistently outperforms the baseline across all datasets. One plausible explanation for this is that representation alignment encourages local representations to stay close to global representations during local training. This in turn helps preserve the knowledge acquired from local data during the server-side calibration process. Figure 4 suggests that instead of acting independently, server-side calibration and representation alignment are complementary for improving the performance of layer-wise training.
5.4 Batch Size
SSL approaches tend to perform better with larger batch sizes due to a greater number of negative samples Chen et al. (2020). To investigate this, we set the local training batch size to 64, 128, 256, 512, and 1024. Figure 5a indicates that layer-wise training approaches, LW-FedMoCoV3 and LW-FedSSL, also tend to improve with increased batch size, similar to end-to-end SSL training. Figure 5b reports peak memory requirements for each setting. Given that Prog-FedSSL does not freeze the layers, its peak memory requirement (at the last stage) is equivalent to that of FedMoCoV3. Memory requirements for FedMoCoV3 and Prog-FedSSL sharply rise as the batch size grows, while those of LW-FedMoCoV3 and LW-FedSSL remain relatively flat.
(a) Linear evaluation of LW-FedMoCoV3.
(b) Linear evaluation of LW-FedSSL.
(c) Linear evaluation of Prog-FedSSL.
Figure 6: Impact of different learning rate strategies.
Figure 7: Linear evaluation results on the STL-10 dataset. (a) Impact of different levels of data heterogeneity. (b) Impact of different numbers of clients.
Figure 8: Learning rate strategies.
5.5 Data Heterogeneity
In this experiment, we use the Tiny ImageNet dataset to construct the local data of clients. We control the data heterogeneity based on data labels and the Dirichlet distribution Ferguson (1973). The concentration parameter $\beta$ in the Dirichlet distribution determines the strength of data heterogeneity, with a lower $\beta$ value indicating a higher degree of heterogeneity. The labels are used here to set up the clients’ data partitions and are not used during training (except for the supervised baseline). Figure 7a presents the linear evaluation results for different $\beta$ values. For the supervised baseline, an end-to-end classification model is trained with vanilla FL using labeled data, and then only the encoder part is kept for linear evaluation. Notably, performance of the supervised approach diminishes as the degree of data heterogeneity increases. In contrast, other approaches employing SSL are relatively more robust to different $\beta$ values. This observation conforms with prior studies Wang et al. (2023) that highlight the robustness of SSL to data heterogeneity in FL environments. Figure 7a indicates that layer-wise training also exhibits robustness to data heterogeneity inherent in end-to-end SSL training.
5.6 Number of Clients
To investigate the impact of the number of clients $N$, we conduct experiments with $N$ set to 5, 10, 30, and 45. For each setting, we uniformly partition the unlabeled data portion of the STL-10 dataset into $N$ local datasets. Therefore, the total number of data samples remains constant across all settings. Linear evaluation results with varying $N$ values are shown in Figure 7b. We can observe that all approaches, except LW-FedSSL, experience a drop in performance as the number of clients increases. Integration of representation alignment mechanism enables LW-FedSSL to maintain good performance even in scenarios with a large number of clients Li et al. (2021).
5.7 Learning Rate Strategy
By default, we use cosine decay scheduling for the learning rate. We explore different learning rate strategies for LW-FedMoCoV3, LW-FedSSL, and Prog-FedSSL, including a fixed learning rate and cosine decay scheduling for each training stage, which we refer to as “cyclic”. Figure 8 illustrates how the learning rate changes at each local epoch for different strategies, and the results are presented in Figure 6. For LW-FedMoCoV3, using a fixed learning rate achieves the best performance, and also for both LW-FedSSL and Prog-FedSSL, simply using a fixed learning rate can be a decent choice. A fixed learning rate may allow layers in later stages to receive effective updates, whereas in cosine decay scheduling, later layers undergo minimal updates due to a small learning rate. While the cyclic learning rate pattern proves to work decently for Prog-FedSSL, it may not be suitable for LW-FedMoCoV3 and LW-FedSSL. As seen in Figure 8, the learning rate for cyclic strategy drops quickly in each stage when only a limited number of rounds are available. This can impact layer-wise training, where a layer receives training only in its corresponding stage. (For LW-FedSSL with a fixed learning rate, we do not use weight transfer, as it demonstrates improved performance.)
6 Conclusion
Edge devices in distributed environments struggle to meet the computation and communication requirements of FedSSL. In response to these challenges, we propose a layer-wise training approach called LW-FedSSL, enabling FL clients to reduce their resource requirements. Our experiments demonstrate that LW-FedSSL has a $\times3$ lower memory requirement, 2.1×2.1$\times$2.1 × lower total FLOPs consumption, and 3.2×3.2$\times$3.2 × cheaper total transmission cost (download + upload), while achieving comparable performance to its end-to-end training counterpart. Furthermore, we explore a progressive training approach called Prog-FedSSL, which remains more resource-efficient than end-to-end training, with significantly improved performance.