### v3

# LW-FedSSL: Resource-efficient Layer-wise Federated Self-Supervised Learning

Abstract
Many studies integrate federated learning (FL) with self-supervised learning (SSL) to take advantage of raw data distributed across edge devices. However, edge devices often struggle with high computation and communication costs imposed by SSL and FL algorithms. To tackle this hindrance, we propose LW-FedSSL, a layer-wise federated self-supervised learning approach that allows edge devices to incrementally train a single layer of the model at a time. We introduce server-side calibration and representation alignment mechanisms to ensure LW-FedSSL delivers performance on par with conventional federated self-supervised learning (FedSSL) while significantly lowering resource demands. In a pure layer-wise training scheme, training one layer at a time may limit effective interaction between different layers of the model. The server-side calibration mechanism takes advantage of the resource-rich FL server to ensure smooth collaboration between different layers of the global model. During local training, the representation alignment mechanism encourages closeness between representations of local models and those of the global model, thereby preserving the layer cohesion established by server-side calibration. With the proposed mechanisms, LW-FedSSL achieves a $\times3$ reduction in memory usage, 2.1×2.1$\times$2.1 × fewer computational operations (FLOPs), and a 3.2×3.2$\times$3.2 × lower communication cost while maintaining the same level of performance as its end-to-end training counterpart. Additionally, we explore a progressive training strategy called Prog-FedSSL, which matches end-to-end training in memory requirements but offers a 1.8×1.8$\times$1.8 × reduction in FLOPs and communication costs. Although Prog-FedSSL is not as resource-efficient as LW-FedSSL, its performance improvements make it a suitable candidate for FL environments with more lenient resource constraints.
keywords:
federated learning , self-supervised learning , layer-wise training , resource-efficient.
journal: Information Fusion
[…]
[label1]organization=Department of Computer Science and Engineering, Kyung Hee University, city=Yongin-si, state=Gyeonggi-do 17104, country=South Korea
[…]
[label2]organization=Vietnam - Korea University of Information and Communication Technology, city=Danang, country=Vietnam
1 Introduction
A significant portion of real-world data, valuable for practical AI applications, resides on edge devices. Such data is often unlabeled, making the adoption of self-supervised learning (SSL) strategies [1, 2, 3] essential. SSL enables models to learn from raw data by generating their own supervisory signals, eliminating the need for labels. Most SSL methods follow centralized training schemes, often requiring an exhaustive data collection process to build a centrally stored dataset. Although some data, such as natural images, can be conveniently collected, it can raise significant concerns for privacy-sensitive data, such as medical data. Meanwhile, decentralized learning approaches like federated learning (FL) [4, 5] enable privacy-preserving collaborative model training. An FL system functions by collecting trained model parameters from edge devices (a.k.a., clients) rather than data. This mechanism safeguards the confidentiality of private data held by different parties. As a result, federated self-supervised learning (FedSSL) approaches have recently emerged as a potential strategy for leveraging raw data in distributed environments [6, 7, 8].
Self-supervised learning (SSL), especially when involving state-of-the-art models like transformers [9, 10], can impose substantial computation demands. Unfortunately, clients in an FL environment often operate with limited computation and communication resources. Many clients may lack the necessary resources to participate in a conventional end-to-end FedSSL process. As a result, the data from these clients is excluded from the training process, potentially degrading the performance of the final model. Therefore, it is crucial to explore resource-efficient SSL approaches for FL systems.
A straightforward way to reduce computation demands on low-memory devices is to conduct the training with a smaller batch size. However, using a small batch size for SSL approaches can diminish the quality of learned representations and compromise performance [1, 7]. Moreover, simply reducing the batch size does not lower the FL communication cost since clients still need to exchange the full model with the server. On the other hand, using a large batch size would exclude low-memory devices from participating in the FL process.
Table 1: FedMoCo (end-to-end) vs. FedMoCo-LW (layer-wise). We use a ViT-Tiny backbone with MoCo v3 as the SSL approach. The FL process involves 10 clients using a batch size of 1024 and the number of communication rounds set to 180. The unlabeled portion of the STL-10 dataset is used to setup the clients’ data, while the labeled training and testing portions are used for linear evaluation. We compare resource requirements for a single client.
Memory
FLOPs
(×$\times 10^{10}$)
Comm.
Accuracy
FedMoCo 15601 27.42 7352 54.99
FedMoCo-LW 3927 9.58 613 46.41
One promising solution is the layer-wise training approach, where a model is systematically trained one layer at a time, gradually increasing its depth. The term “layer” can refer to either an individual layer or a block comprising multiple layers. In layer-wise training, only a single layer is active at a time for training, significantly reducing computation requirements on FL clients. Moreover, only the active layer needs to be exchanged with the server, further reducing communication costs. Table 1 shows a significant reduction in a client’s resource requirements for layer-wise training compared to end-to-end training.
Although layer-wise training significantly reduces resource requirements, it may not be able to match the performance of end-to-end training, as shown in Table 1. In each stage of layer-wise training, a new layer is sequentially attached to the model, and all preceding layers are frozen, excluding them from further updates. Given a fixed number of FL rounds for each stage, it can be challenging to train the newly added layer to work seamlessly with the preceding layers, especially if the number of rounds is small. Although increasing the number of rounds could potentially alleviate this problem, it would impose extra communication and computation overhead on the clients, canceling the benefits of layer-wise training.
To address these challenges, we propose LW-FedSSL by introducing two simple, yet effective mechanisms: server-side calibration and representation alignment, designed to complement the layer-wise training process in an FL paradigm. Although clients in an FL environment may have limited resources, the central server typically possesses significant computation power. The server-side calibration mechanism leverages these resources to train the global model in an end-to-end manner, establishing layer cohesion. This mechanism solely operates on the central server and incurs no additional overhead on the clients. To preserve layer cohesion during local training, the representation alignment mechanism encourages alignment between local models and the server-side-trained global model.
Our contributions can be summarized as follows:
  • 1.
    We conduct an in-depth exploration of layer-wise training for SSL within the FL paradigm, considering various empirical aspects.
  • 2.
    We introduce LW-FedSSL, a novel layer-wise training framework designed to improve the resource efficiency of FL clients. By leveraging server-side calibration and representation alignment mechanisms, LW-FedSSL can compete with conventional end-to-end training counterpart while significantly lowering resource requirements.
  • 3.
    We explore Prog-FedSSL, a progressive training strategy. While this approach demands more resources than the layer-wise approach, it has the potential to enhance performance. Given its lower resource efficiency compared to LW-FedSSL, Prog-FedSSL is better suited for FL environments with more flexible resource constraints.
  • 4.
    We conduct a thorough evaluation of our proposed approaches through various FL settings and benchmarks.
2 Background
In this section, we briefly introduce fundamental concepts related to federated learning and self-supervised learning.
Figure 1: Four key steps in an FL communication round. (i) The server distributes a base global model $M$ to all participating clients. (ii) Each client $i\in[1,N]$ performs training on its local dataset $D^{i}$ to produce a local model $M^{i}$. (iii) Then, the local models are transmitted back to the server. (iv) The sever finally aggregates the received local models through weighted averaging to update the global model: $M\leftarrow\sum^{N}_{i=1}w^{i}M^{i}$ [5].
2.1 Federated Learning
A typical FL process involves a central coordinating server and a set of client devices, each of which stores private data [5]. The goal of FL is to train a global model $M$, by learning from the data residing on clients while ensuring their privacy. The FL objective can be expressed as
$M^{*}=\operatorname*{arg\,min}_{M}\sum_{i=1}^{N}w^{i}\mathcal{L}(M,D^{i}),$ (1)
where $N$ is the number of clients and $\mathcal{L}$ is the local loss. Here, $w^{i}=|D^{i}|/|D|$, which is the weight term of the $i$-th client, with $|D^{i}|$ denoting the number of samples in the local dataset $D^{i}$ and $|D|$ denoting the total number of samples in ${D=\bigcup_{i=1}^{N}D^{i}}$. As shown in Fig. 1, a single FL communication round consists of four main steps, which are repeated for a number of rounds $R$.
2.2 Self-Supervised Learning
Self-supervised learning (SSL) offers a means to utilize unlabeled data, which is often more readily available than labeled data for model training. SSL generates its own supervisory signals to learn valuable representations from the unlabeled data. We select MoCo v3 [11] as the SSL approach in our work. As shown in Fig 2, MoCo v3 features a Siamese network structure with an actively trained online branch, and a target branch, which is a moving-averaged copy of the online branch. The online branch consists of an encoder $F$, a projection head $H$, and a prediction head $P$. Meanwhile, the target branch consists of a momentum encoder $F_{k}$ and a momentum projection head $H_{k}$. Given an input sample $x$, it undergoes augmentation, creating two views, $x_{q}$ and $x_{k^{+}}$. These views are then fed into the online and target branches, respectively, to obtain a vector $q$ and its corresponding positive pair, $k^{+}$. Likewise, negative pairs, $k^{-}$, are derived from other samples within the same batch through the target branch. The online branch is trained using the InfoNCE loss [12],
$\ell(q,k,\tau)=-\log\frac{\exp(q\cdot k^{+}/\tau)}{\exp(q\cdot k^{+}/\tau)+\sum_{j=1}^{K}\exp(q\cdot k_{j}^{-}/\tau)},$ (2)
where $k=(k^{+},\{k_{j}^{-}\}_{j=1}^{K})$ and $\tau$ is the temperature parameter. Using a momentum parameter $\mu$, the target branch is updated as: $F_{k}\leftarrow\mu F_{k}+(1-\mu)F$ and $H_{k}\leftarrow\mu H_{k}+(1-\mu)H$.
2.3 Federated Self-Supervised Learning
Although numerous studies on FL primarily focus on supervised learning, this means that device owners are tasked with labeling data, which is a tedious process. Consequently, client devices often lack labeled data due to limitations in data collection methods, labeling constraints, and insufficient annotation expertise and resources. In response, federated self-supervised learning (FedSSL) emerges as a solution to extend the applicability of FL to new domains where labeled data is scarce or difficult to acquire [13, 14, 15]. Many studies on FedSSL build on state-of-the-art SSL approaches like SimCLR [1], MoCo [2], and BYOL [16]. These studies adapt SSL techniques to address specific FL tasks [17, 18] and challenges at hand [19, 20, 21, 6]. In the context of FedSSL, the local dataset $D^{i}$ within a client $i$ is unlabeled, and thus, the local training process involves utilizing a form of SSL loss to train the local model $M^{i}$ on $D^{i}$. When using MoCo v3 as the SSL approach, this local model can comprise an encoder $F^{i}$ and its accompanying MLP heads, $H^{i}$ and $P^{i}$.
3 Related Work
Figure 2: Self-supervised learning with MoCo v3 [11]. We omit the negative samples for clarity.
Layer-wise training was initially introduced in the context of deep belief networks and restricted Boltzmann machines [22, 23]. The original motivation behind its development was to address the challenges associated with training deep neural networks, such as vanishing gradients. While layer-wise training has become less common over the years, it can still be effectively employed as a resource-efficient strategy within FL systems. Training the model incrementally, one layer at a time, offers a distinct resource-saving advantage over end-to-end training. To date, only a very limited number of studies [24, 25] have explored layer-wise training in the FL context, and even fewer [25] have examined its application in the domain of FedSSL.
The study in [26] investigates layer-wise SSL in the speech domain. While it was intended for an on-device training scenario, the authors mainly examined it in a centralized setting. In [25], the authors propose an approach called federated layer-wise learning (FLL), based on SimCLR [1]. They introduce a depth dropout technique that randomly drops frozen layers during training to further reduce resource overhead. However, FLL is only examined with a large number of communication rounds in each layer-wise training stage (ranging from 4k to 12k), potentially placing significant strain on clients. Additionally, the exploration of FLL is limited in scope, missing comprehensive evaluation across diverse FL settings and benchmark datasets. In contrast to the above studies, we aim to offer a more in-depth exploration of layer-wise training in the context of FedSSL. Moreover, as we have previously demonstrated, layer-wise training alone may not achieve the same level of performance as end-to-end training. To bridge this gap, we introduce LW-FedSSL, a layer-wise training approach that can compete end-to-end training with significantly lower client resource requirements. One study [24] investigates the progressive training approach for FL known as ProgFed. Although similar to layer-wise training in gradually increasing the model depth, ProgFed differs by training all existing layers at each stage. However, ProgFed primarily focuses on federated supervised learning tasks that require labeled data for training. The nature of a progressive training approach within the FedSSL paradigm remains unexplored. To fill this gap, we introduce a progressive training strategy for FedSSL, referred to as Prog-FedSSL.
Figure 3: Layer-wise training process across different stages for the $i$-th client. Each stage $s\in[1,S]$ increases the depth of the encoder $F$ by adding a new layer $L_{s}$. During stage $s$, only the corresponding layer $L_{s}$ is actively updated, while prior layers (i.e., $L_{1}$ to $L_{s-1}$) are frozen.
4 Proposed Method
4.1 FedMoCo-LW
Firstly, we discuss the pure layer-wise training baseline, referred to as FedMoCo-LW. Fig. 3 illustrates the local training process of an FL client across different stages. For an encoder $F$ with a total of $S$ layers, the training process is divided into $S$ stages. Each stage $s\in[1,S]$ runs for a fixed number of FL communication rounds, $R_{s}$. The training process starts with an empty encoder $F$ without any layers, and each stage $s$ sequentially adds a new layer $L_{s}$ to $F$, gradually increasing its depth. During stage $s$, only the corresponding layer $L_{s}$ is updated, while all prior layers $L_{1}$ to $L_{s-1}$ within $F$ are frozen, serving only for inference. This strategy saves memory and computation, benefiting resource-constrained devices in FL environments. Moreover, in each communication round, clients are only required to send the active layer $L_{s}$ (and the MLP heads) to the server for aggregation. Likewise, the server only needs to broadcast the aggregated layer $L_{s}$ back to clients. This reduces both upload and download communication costs for clients. Although layer-wise training offers better resource efficiency than end-to-end training, there remains a notable performance gap in linear evaluation, as we have shown in Table 1.
Figure 4: LW-FedSSL at stage $s$. The server-side calibration mechanism (top) trains all existing layers in the current sub-model. The representation alignment mechanism (bottom) aligns the representations of the local model with those of the global model during the local training process.
4.2 LW-FedSSL
To address the performance gap between layer-wise and end-to-end training for FedSSL, we present LW-FedSSL by introducing two crucial mechanisms: server-side calibration, integrated at the server, and representation alignment, integrated into the local training process. Fig. 4 illustrates these two mechanisms within LW-FedSSL at stage $s$.
Server-side calibration. In layer-wise training, only the most recently added layer, $L_{s}$, is actively trained by clients at any given stage $s$. This may limit prior layers’ ability to adapt to the newly added layer, potentially degrading the model’s performance. Given a limited number of rounds $R_{s}$ available for each stage, solely training $L_{s}$ may prove insufficient to match the training received by the end-to-end strategy. On the other hand, increasing the number of rounds could introduce additional resource demands on clients. To tackle this, we can take advantage of resources available at the FL server. The server can contribute by training the current sub-model on a small auxiliary dataset, $D^{g}$, at the end of every communication round. Specifically, following the aggregation process, the server can train the aggregated encoder $F$ across all existing layers (i.e., $L_{1}$ to $L_{s}$) before broadcasting $F$ back to the clients. The auxiliary dataset $D^{g}$ may originate from a public dataset with a different data distribution than that of the clients’ data, as the server cannot directly access data within clients. Furthermore, $D^{g}$ may typically be unlabeled, requiring server-side training to also utilize an SSL approach. Training across all existing layers in the current sub-model plays a crucial role in calibrating the layers to function harmoniously. Also, our assumption regarding the auxiliary data at the server can be reasonably met since, in practice, the server may require a small dataset for various purposes, such as performance monitoring, pre-training [27, 7, 28], distillation [27], and other tasks in the FL context [29, 19, 30]. Particularly for FedSSL tasks, probing datasets may be necessary at the server to monitor and ensure smooth training progress.
Algorithm 1 LW-FedSSL (Server-side)
Input: encoder $F$, projection head $H$, prediction head $P$, auxiliary dataset $D^{g}$, number of epochs $E$, momentum $\mu$, temperature $\tau$, number of stages $S$, number of rounds per stage $R_{s}$, ${R\leftarrow\sum_{s=1}^{S}R_{s}}$
Output: encoder $F$
1:Server executes:
2:Initialize stage: $s=0$
3:for round $r=1,2,\dots,R$ do
4:     if $r>\sum_{j=1}^{s}R_{j}$ then
5:         $s\leftarrow s+1$
6:         Append a new layer $L_{s}$ to $F$
7:     Unfreeze $F$
8:     $F,H,P\leftarrow\text{Server\_SSL}(D^{g},E,\mu,\tau,F,H,P)$
9:     Freeze $F$ up to layer $L_{s-1}$
10:     for client $i=1,2,\dots,N$ in parallel do
11:         $L_{s}^{i},H^{i},P^{i}\leftarrow\text{Local\_SSL}(i,F,H,P)$triangleright Algo. 2
12:         $w^{i}=\frac{|D^{i}|}{|D|}$, where $D=\bigcup_{i=1}^{N}D^{i}$
13:     $L_{s},H,P\leftarrow\text{Aggregation}(\{L^{i}_{s},H^{i},P^{i},w^{i}\}^{N}_{i=1})$
14:     Update $L_{s}$ in $F$
15:return $F$
Representation alignment. FL clients often exhibit distinct local data distributions resulting from diverse device usage patterns and application settings. Such data heterogeneity can introduce bias to the local training process, negatively affecting the performance of the global model. To mitigate this, some studies [31, 32] encourage local training processes across different clients to converge toward parameters that are closely aligned with the global parameters. Often, this is achieved by aligning the representations of local models with those of the global model while simultaneously pushing them away from old representations (from old local models of previous FL rounds) [31, 33]. In contrast, our representation alignment mechanism does not require old local parameters. As a result, clients no longer need to store and perform inference on old parameters, reducing the overhead. Given a positive pair of views, $x_{q}$ and $x_{k^{+}}$, our approach uses local encoder $F^{i}$ and the global encoder $F$ to extract local representation $z^{i}=F^{i}(x_{q})$ and positive global representation $z^{+}=F(x_{k^{+}})$. Negative global representations $z^{-}$ are extracted from other samples. Similar to Equation 2, we can calculate the alignment loss as
$\ell(z^{i},z,\tau)=-\log\frac{\exp(z^{i}\cdot z^{+}/\tau)}{\exp(z^{i}\cdot z^{+}/\tau)+\sum_{j=1}^{K}\exp(z^{i}\cdot z_{j}^{-}/\tau)},$ (3)
where $z=(z^{+},\{z_{j}^{-}\}^{K}_{j=1})$. This loss can be integrated into local training as an auxiliary loss for regularization, with a weight term $\alpha$ to control its contribution.
The server-side calibration mechanism trains all existing layers in the global model to facilitate smooth collaboration between them. However, during local training, updates only to the active layer may disrupt this cohesion. To mitigate this, representation alignment encourages the updates of the local model to remain closely aligned with the parameters of the server-side trained global model. This helps alleviate complete overwriting and preserves the layer cohesion established by server-side calibration, while also allowing local models to learn new information from their local datasets. Additionally, representation alignment prevents extensive divergence between local models of different clients, enabling a more effective FL aggregation process.
In Section 5.3, we empirically show that server-side calibration and representation alignment are complementary for improving the effectiveness of layer-wise training. Integrating these mechanisms enables LW-FedSSL to compete with end-to-end training, all the while substantially reducing resource demands on FL clients. The details of LW-FedSSL are presented in Algorithms 1 and 2. Algorithm 1 outlines the server-side execution, which handles the gradual increase in the model depth and training stage, the server-side calibration mechanism, and the aggregation process. Each client executes Algorithm 2, performing local SSL training with representation alignment as regularization.
Algorithm 2 LW-FedSSL (Client-side)
Input: local dataset $D^{i}$, number of local epochs $E$, momentum $\mu$, temperature $\tau$, weight term $\alpha$
Output: layer $L^{i}_{s}$, local projection head $H^{i}$, local prediction head $P^{i}$
1:Client executes: Local_SSL($i,F,H,P$):
2:Online branch: $F^{i}\leftarrow F$ and $H^{i}\leftarrow H$ and $P^{i}\leftarrow P$
3:Target branch: $F^{i}_{k}\leftarrow F$ and $H^{i}_{k}\leftarrow H$
4:for epoch $e=1,2,\dots,E$ do
5:     for each batch $x\in D^{i}$ do
6:         $x_{1}\leftarrow\text{{Augment}}(x)$ and $x_{2}\leftarrow\text{{Augment}}(x)$
7:         $z_{1}\leftarrow F(x_{1})$ and $z_{2}\leftarrow F(x_{2})$
8:         $z_{1}^{i}\leftarrow F^{i}(x_{1})$ and $z_{2}^{i}\leftarrow F^{i}(x_{2})$
9:         $q_{1}^{i}\leftarrow P^{i}(H^{i}(z_{1}^{i}))$ and $q_{2}^{i}\leftarrow P^{i}(H^{i}(z_{2}^{i}))$
10:         $k_{1}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{1}))$ and $k_{2}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{2}))$
11:         $\ell_{\text{con}}\leftarrow\ell(q^{i}_{1},k^{i}_{2},\tau)+\ell(q^{i}_{2},k^{i}_{1},\tau)$triangleright Eq. (2)
12:         $\ell_{\text{align}}\leftarrow\ell(z^{i}_{1},z_{2},\tau)+\ell(z^{i}_{2},z_{1},\tau)$triangleright Eq. (3)
13:         $\mathcal{L}\leftarrow\ell_{\text{con}}+\alpha\ell_{\text{align}}$
14:         $F^{i},H^{i},P^{i}\leftarrow\text{Update}(\nabla\mathcal{L})$
15:         $F^{i}_{k},H^{i}_{k}\leftarrow\text{Momentum\_Update}(F^{i},H^{i},\mu)$
16:$L_{s}^{i}\leftarrow\text{Get active layer from }F^{i}$
17:return $L_{s}^{i},H^{i},P^{i}$
Table 2: Given an encoder $F$ with a total of $S$ layers, we compare the characteristics of different training approaches at stage $s\in[1,S]$.
At Stage $\mathbf{s}$
Training
Number of
Stages
Initial
Layers
New
Layer
Frozen
Layers
Active
Layers
Server to Client
Client to Sever
FedMoCo end-to-end $L_{1}$ to $L_{S}$ $L_{1}$ to $L_{S}$ $L_{1}$ to $L_{S}$ $L_{1}$ to $L_{S}$
FedMoCo-LW layer-wise $S$ $L_{s}$ $L_{1}$ to $L_{s-1}$ $L_{s}$ $L_{s}$ $L_{s}$
LW-FedSSL layer-wise $S$ $L_{s}$ $L_{1}$ to $L_{s-1}$ $L_{s}$ $L_{1}$ to $L_{s}$ $L_{s}$
Prog-FedSSL progressive $S$ $L_{s}$ $L_{1}$ to $L_{s}$ $L_{1}$ to $L_{s}$ $L_{1}$ to $L_{s}$
4.3 Prog-FedSSL
Inspired by ProgFed [24], we explore progressive training for FedSSL, denoted as Prog-FedSSL. Similar to layer-wise training, progressive training also starts with an empty encoder $F$, appending a new layer $L_{s}$ to $F$ at each stage $s$. However, in contrast to updating only a single layer $L_{s}$, progressive training updates all existing layers (i.e., $L_{1}$ to $L_{s}$) during each stage $s$. This distinction results in higher computation and communication costs for Prog-FedSSL than LW-FedSSL. Nonetheless, when compared with end-to-end training, which requires clients to train a full model with $S$ layers (i.e., $L_{1}$ to $L_{S}$) at each round, Prog-FedSSL enables clients to train only a sub-model with $s$ layers $(s\leq S)$. Moreover, clients only need to exchange the sub-model with the server. In essence, Prog-FedSSL saves both computation and communication costs for $(S-s)$ layers at each stage $s$ compared to end-to-end training. As we will demonstrate in Section 5.2, progressive training has the potential to outperform both layer-wise and end-to-end training approaches.
Nevertheless, it is important to recognize that the resource requirement of progressive training gradually increases as the stages progress, becoming equivalent to that of end-to-end training at the final stage (i.e., when $s=S$). This growing demand could prevent resource-constrained devices from participating in the training, raising similar limitations faced by end-to-end approach, leading to the loss of valuable data from these devices. Meanwhile, the resource requirement of layer-wise training can be significantly lower at any stage due to training a single layer $L_{s}$ at a time. We compare the different characteristics of FedMoCo (end-to-end), FedMoCo-LW (layer-wise), LW-FedSSL, and Prog-FedSSL in Table 2. In A, Fig. 15 illustrates FedMoCo, FedMoCo-LW, and Prog-FedSSL at stage $s$.
5 Experiment
5.1 Experimental setup
Unless otherwise stated, we use the following settings in our experiments.
Model. We use a ViT-Tiny [34] backbone with 12 transformer blocks (i.e., $S=12$) as encoder $F$. We use an input dimension of $32\times 32\times 3$ with a patch size of 4. The projection head $H$ and the prediction head $P$ are implemented as a 3-layer MLP and a 2-layer MLP, respectively. For both $H$ and $P$, the hidden layer dimensions are set to 4096 and the output dimension is set to 256.
[…]
Evaluation. Following common practices, we primarily use linear evaluation for performance comparison [37, 38, 1]. The projection head $H$ and prediction head $P$ are discarded, and a linear classifier is trained on top of the frozen encoder $F$ for 20 epochs using a batch size of 256. We use the AdamW optimizer with a base learning rate of 3e-2 and a weight decay of 1e-5. The learning rate decays with a cosine schedule. We use the RandAugment function available in the PyTorch [39] framework for augmentation. For fine-tuning, we set the base learning rate to 1e-3 and the number of epochs to 40 with a warmup period of 10 epochs. Other settings are similar to those of linear evaluation.
Datasets. We uniformly split the unlabeled data portion of the STL-10 dataset [40] to construct the local data of FL clients. To evaluate the performance, we use the CIFAR-10/100 [41], STL-10 (labeled portion), and Tiny ImageNet [42] datasets. For LW-FedSSL, we use only 10% of CIFAR-10 training data, excluding labels, as the auxiliary dataset $D^{g}$.
Baselines. In line with existing FedSSL studies [13, 20, 6], we establish end-to-end training baselines by integrating state-of-the-art SSL algorithms—MoCo v3 [11], BYOL [16], and SimCLR [1]—with FL [5] to create FedMoCo, FedBYOL, and FedSimCLR, respectively. We include FedMoCo-LW and FLL [25] with depth dropout (DD) as layer-wise training baselines. We set the depth dropout rate to 50%, as in [25]. For FedSimCLR and FLL, we use a batch size of 512. We also include MoCo v3 as the centralized upper bound. We mainly compare LW-FedSSL against FedMoCo and FedMoCo-LW, given that they are all based on the same SSL approach.
Table 3: Linear evaluation and fine-tuning performance of different approaches. We conduct the experiments three times with different seeds. We compare max memory, total FLOPs, and communication costs (download + upload) for a client. Three best values, except for the centralized MoCo v3, are marked in bold.
Cost Linear Evaluation Fine-tune
Memory FLOPs Comm. CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
MoCo v3 70.82 $\pm$ 0.06 46.09 $\pm$ 0.45 63.84 $\pm$ 0.42 28.61 $\pm$ 0.45 86.27 $\pm$ 0.26 64.57 $\pm$ 0.38 69.78 $\pm$ 0.91 43.95 $\pm$ 0.48
Scratch 23.47 $\pm$ 1.20 5.83 $\pm$ 0.42 17.50 $\pm$ 1.00 2.68 $\pm$ 1.09 63.23 $\pm$ 0.94 41.69 $\pm$ 0.58 44.08 $\pm$ 0.13 27.99 $\pm$ 0.33
FedMoCo 1.00×1.$0$ 1.00×1.$0$ 1.00×1.$0$ 60.64 $\pm$ 0.14 34.66 $\pm$ 0.42 54.96 $\pm$ 0.03 20.55 $\pm$ 0.28 79.46 $\pm$ 0.32 58.70 $\pm$ 0.32 62.59 $\pm$ 0.36 40.62 $\pm$ 0.28
FedBYOL 1.00×1.$0$ 1.00×1.$0$ 1.00×1.$0$ 55.38 $\pm$ 0.03 28.64 $\pm$ 0.04 51.01 $\pm$ 0.07 16.34 $\pm$ 0.05 77.25 $\pm$ 0.11 56.92 $\pm$ 0.39 60.36 $\pm$ 0.18 39.58 $\pm$ 0.39
FedSimCLR 0.99×0.$9$ 1.49×1.49$\times$1.49 × 1.00×1.$0$ 62.59 $\pm$ 0.03 36.94 $\pm$ 0.08 57.89 $\pm$ 0.06 21.81 $\pm$ 0.07 79.50 $\pm$ 0.39 58.45 $\pm$ 0.15 63.15 $\pm$ 0.24 41.43 $\pm$ 0.27
[…]
Prog-FedSSL 1.00×1.$0$ 0.57×0.57$\times$0.57 × 0.54×0.54$\times$0.54 × 65.76 $\pm$ 0.10 39.45 $\pm$ 0.06 60.83 $\pm$ 0.07 23.02 $\pm$ 0.51 86.21 $\pm$ 0.23 64.01 $\pm$ 0.41 69.45 $\pm$ 0.54 41.04 $\pm$ 0.13
(a) Memory.
(b) FLOPs.
(c) Download.
(d) Upload.
Figure 5: Computation and communication resources required for a local training process. (a) Memory requirement. (b) FLOPSs consumption. (c) Download size. (d) Upload size.
(a) Linear evaluation.
(b) Peak memory requirement.
Figure 6: Impact of batch size on different approaches. (a) Linear evaluation results on the STL-10 dataset, and (b) peak memory requirements for different batch sizes.
5.2 Performance and Resource Comparison
In Table 3, we compare the performance of different approaches. In terms of linear evaluation, LW-FedSSL demonstrates comparable performance to the end-to-end training of FedMoCo and significantly outperforms the layer-wise baseline FedMoCo-LW. LW-FedSSL also outperforms or holds a competitive position against FedMoCo in fine-tuning. The progressive training approach, Prog-FedSSL, outperforms all other approaches in both linear evaluation and fine-tuning performance. Interestingly, the fine-tuning performance of FedMoCo is even lower than that of FedMoCo-LW in some cases. Please note that we apply the same fine-tuning settings, which may not be optimal for all approaches.
In Fig. 5, we present resource requirements for a single client. Fig. 5a measures the memory usage at each local epoch, showing that the memory requirement of LW-FedSSL is $\times3$ lower than that of FedMoCo at peak usage. The memory usage of Prog-FedSSL gradually increases and eventually matches that of FedMoCo in the final stage. Fig. 5b shows that LW-FedSSL also maintains a lower FLOPs requirement than FedMoCo, with 2.1×2.1$\times$2.1 × lower total FLOPs consumption. Additional memory and FLOPs requirements of LW-FedSSL over FedMoCo-LW arise from the representation alignment mechanism, which requires inference on the global model to obtain global representations.
In Figures 5c and 5d, we investigate the communication cost required for model exchange (download and upload) between the server and a client. Note that we only consider the encoder part for comparison, excluding the projection and prediction heads, which are common to all approaches with constant communication costs. In FedMoCo, a client needs to exchange the full encoder with the server at each communication round. Meanwhile, in FedMoCo-LW, a client only needs to exchange the active layer (i.e., $L_{s}$). Thus, both FedMoCo and FedMoCo-LW result in constant download and upload costs across different communication rounds. For Prog-FedSSL, both download and upload costs gradually increase as the number of active layers (i.e., $L_{1}$ to $L_{s}$) grows with the stage $s$. As for LW-FedSSL, a client needs to download layers $L_{1}$ to $L_{s}$ from the server, as these are updated through the server-side calibration mechanism. However, the client uploads a single layer $L_{s}$, as only it is updated by the local training process. Therefore, the download cost of LW-FedSSL increases with each stage $s$, while the upload cost remains constant. Upon comparing the total communication cost over 180 rounds, LW-FedSSL is 1.8×1.8$\times$1.8 × cheaper in terms of download cost and 12×12$\times$12 × cheaper in terms of upload cost compared to FedMoCo.
It is important to recognize that we are considering the best-case scenario for the baseline FedMoCo-LW (and FLL+DD in Table 3), where each client participates in every training round. In this case, only the aggregated layer $L_{s}$ needs to be downloaded from the server, resulting in lower communication costs than LW-FedSSL. However, when new clients join or existing clients miss certain training stages, they need to download the current submodel ($L_{1}$ to $L_{s}$) from the server, similar to LW-FedSSL.
5.3 Ablation Study
Figure 7: Linear evaluation performance of FedMoCo-LW, LW-FedSSL, and its two variants, each incorporating either server-side calibration or representation alignment. Weight transfer between stages is not performed for these two variants, as they demonstrate better performance without it.
In Fig. 7, we compare the baseline FedMoCo-LW with LW-FedSSL and its two variants, which solely incorporate either server-side calibration or representation alignment. Incorporating server-side calibration improves performance on the CIFAR-10 dataset (from which the auxiliary data is sampled), but leads to a decline on other datasets. This decline may result from conflicting update directions between training on auxiliary data and the local datasets. Compared to the baseline, incorporating representation alignment obtains better performance on the STL-10 dataset, with comparable performance on the others. Meanwhile, LW-FedSSL, which incorporates both mechanisms, consistently outperforms the baseline across all datasets. The server-side calibration mechanism facilitates smooth collaboration between different layers of the model. Meanwhile, representation alignment helps maintain layer cohesion established by server-side calibration during the local training process. Fig. 7 suggests that instead of acting independently, server-side calibration and representation alignment are complementary for improving the performance of layer-wise training.
Table 4: Linear evaluation results of LW-FedSSL with different amounts of auxiliary data.
Ratio CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
10% ($D^{g}$ only) 43.75 20.03 40.97 9.77
1% 47.80 23.69 43.65 12.76
10% 61.83 32.25 55.15 17.26
100% 79.63 47.09 69.52 26.75
5.4 Amount of Auxiliary Data
LW-FedSSL relies on the auxiliary dataset $D^{g}$ for the server-side calibration mechanism. We vary the amount of auxiliary data at the server by sampling 1%, 10%, and 100% of the CIFAR-10 training data. We also conduct a centralized end-to-end SSL training using only $D^{g}$ from the sever. As expected, the results in Table 4 indicate that the performance of LW-FedSSL improves with an increase in the amount of data.
(a) Linear evaluation of FedMoCo-LW.
(b) Linear evaluation of LW-FedSSL.
(c) Linear evaluation of Prog-FedSSL.
Figure 8: Impact of different learning rate strategies.
5.5 Batch Size
SSL approaches tend to perform better with larger batch sizes due to a greater number of negative samples [1]. To investigate this, we set the local training batch size to 64, 128, 256, 512, and 1024. Fig. 6a indicates that layer-wise training approaches, FedMoCo-LW and LW-FedSSL, also tend to improve with increased batch size, similar to end-to-end SSL training. Fig. 6b reports peak memory requirements for each batch size setting. Given that Prog-FedSSL does not freeze the layers, its peak memory requirement (at the last stage $S$) is equivalent to that of FedMoCo. Memory requirements for FedMoCo and Prog-FedSSL sharply rise as the batch size grows, while those of FedMoCo-LW and LW-FedSSL remain relatively flat. When using comparable memory resources (as seen with the batch size of 64), LW-FedSSL outperforms FedMoCo while also retaining its layer-wise communication efficiency advantage.
Figure 9: Different levels of data heterogeneity. Linear evaluation results are shown for the STL-10 dataset.
Figure 10: Total number of clients. Linear evaluation results are shown for the STL-10 dataset.
5.6 Data Heterogeneity
Data heterogeneity is one of the main challenges in an FL environment. In this experiment, we use the Tiny ImageNet dataset to construct the local data of clients. We control the data heterogeneity based on data labels and the Dirichlet distribution [43]. The concentration parameter $\beta$ in the Dirichlet distribution determines the strength of data heterogeneity, with a lower $\beta$ value indicating a higher degree of heterogeneity. The labels are used here to set up the clients’ data partitions and are not used during training (except for the supervised baseline). Fig. 9 presents the linear evaluation results for different $\beta$ values. For the supervised baseline, an end-to-end classification model is trained with vanilla FL using labeled data, and then only the encoder part is kept for linear evaluation. Notably, performance of the supervised approach diminishes as the degree of data heterogeneity increases. In contrast, other approaches employing SSL are relatively more robust to different $\beta$ values. This observation conforms with prior studies [8] that highlight the robustness of SSL to data heterogeneity in FL environments. Fig. 9 indicates that layer-wise training also exhibits robustness to data heterogeneity inherent in end-to-end SSL training.
Figure 11: Number of participating clients in each FL round. Linear evaluation results are for STL-10.
Figure 12: Learning rate strategies.
5.7 Total Number of Clients
(a) Uniform allocation.
(b) Right-skewed allocation. (high to low)
(c) Left-skewed allocation. (low to high)
Figure 13: Number of rounds per stage.
(a) Linear evaluation of FedMoCo-LW.
(b) Linear evaluation of LW-FedSSL.
(c) Linear evaluation of Prog-FedSSL.
Figure 14: Impact of number of rounds per stage.
To investigate the impact of the number of clients $N$, we conduct experiments with $N$ set to 5, 10, 30, and 45. For each setting, we uniformly partition the unlabeled data portion of the STL-10 dataset into $N$ local datasets. Therefore, the total number of data samples remains constant across all settings. Linear evaluation results with varying $N$ values are shown in Fig. 10. We can observe that all approaches, except LW-FedSSL, experience a drop in performance as the number of clients increases. The representation alignment mechanism helps prevent significant divergence between local models, particularly in scenarios with a large number of clients [31, 33], enabling LW-FedSSL to maintain robust performance.
5.8 Number of Participating Clients
In FL environments, it is common for only a subset of clients to participate in training at any given time. Therefore, we experiment with the number of clients active for training at each FL round. We set the total number of clients to 45 and randomly sample 5, 10, 20, and 45 clients at each FL round. The linear evaluation results in Fig. 11 demonstrate that LW-FedSSL can maintain its performance even with a small proportion of participating clients in each FL round.
5.9 Learning Rate Strategy
By default, we use cosine decay scheduling for the learning rate. We explore different learning rate strategies for FedMoCo-LW, LW-FedSSL, and Prog-FedSSL, including a fixed learning rate and cosine decay scheduling for each training stage, which we refer to as “cyclic”. Fig. 12 illustrates how the learning rate changes at each local epoch for different strategies, and the results are presented in Fig. 8. For FedMoCo-LW, using a fixed learning rate achieves the best performance, and also for both LW-FedSSL and Prog-FedSSL, simply using a fixed learning rate can be a decent choice. A fixed learning rate may allow layers in later stages to receive effective updates, whereas in cosine decay scheduling, later layers undergo minimal updates due to a small learning rate. While the cyclic learning rate pattern proves to work decently for Prog-FedSSL, it may not be suitable for FedMoCo-LW and LW-FedSSL. As seen in Fig. 12, the learning rate for cyclic strategy drops quickly in each stage when only a limited number of rounds are available. This can impact layer-wise training, where a layer receives training only in its corresponding stage.
5.10 Distribution of Number of Rounds Per Stage
In our experiments, the number of rounds for each layer-wise training stage is uniformly distributed. As shown in Fig. 13, we explore two alternative settings: one where a larger number of rounds is allocated to earlier stages, and the other where a larger number of rounds is allocated to later stages. We refer to these settings as “right-skewed allocation” and “left-skewed allocation”, respectively. The total number of rounds across all settings remains constant at 180. The results are shown in Fig. 14.
For FedMoCo-LW, the left-skewed setting results in the worst performance. In this setting, earlier layers receive updates for a small number of rounds, potentially lowering the quality of their output features. Even though later layers are trained for a large number of rounds, poor output features from the early layers may degrade the output of subsequent layers. The other two settings achieve comparable performance. As for LW-FedSSL, it allows early layers to continuously receive updates through the server-side calibration mechanism, which improves the performance of the left-skewed setting in most cases.
In the case of Prog-FedSSL, none of the layers from prior stages are frozen, allowing layer $L_{s}$ to receive training for all the rounds allocated through stages $s$ to $S$. This means that in a left-skewed allocation scenario, if we assign 10 rounds to stage $(S-1)$ and 50 rounds to stage $S$, then the layer $L_{S-1}$ undergoes training for a total of 60 rounds (since it is not frozen), while $L_{S}$ is trained for 50 rounds. However, if we reverse the scenario in the right-skewed allocation, i.e., assign 50 rounds to stage $(S-1)$ and 10 rounds to stage $S$, the layer $L_{S-1}$ still undergoes training for 60 rounds, while $L_{S}$ is trained only for 10 rounds. As a result, the left-skewed setting achieves better performance than the right-skewed setting; however, it would also consume more resources. The same case also applies to the server-side calibration mechanism of LW-FedSSL.
6 Conclusion
Edge devices in distributed environments often struggle to meet the computation and communication requirements of FedSSL. In response to these challenges, we propose a layer-wise training approach named LW-FedSSL, which allows FL clients to perform self-supervised learning with improved resource efficiency. By leveraging server-side calibration and representation alignment mechanisms, LW-FedSSL achieves comparable performance to end-to-end training while significantly reducing resource demands. Specifically, it requires $\times3$ less memory, consumes 2.1×2.1$\times$2.1 × fewer FLOPs, and reduces total transmission costs (download and upload) by 3.2×3.2$\times$3.2 ×. Furthermore, we explore a progressive training approach called Prog-FedSSL. Although it offers lower resource efficiency than LW-FedSSL, it remains more efficient than end-to-end training, while also improving performance.
Appendix A Additional Details
A.1 Details on FLOPs Calculation
Here, we discuss the details on FLOPs calculation used in Section 5.2. In layer-wise training, calculating FLOPs for inactive (frozen) layers involves simply considering FLOPs associated with the forward pass. However, for active layers, the calculation must account for both the forward and backward passes. While numerous works have addressed FLOPs calculation for the forward pass (inference), limited studies are available on the practice of FLOPs calculation for the backward pass. Notably, existing works related to layer-wise training [24, 25] also do not provide specific details on how FLOPs are computed. Some studies [44, 45, 46, 47] suggest that the number of operations in a backward pass of a neural network is often twice that of a forward pass. Following these studies, we adopt a backward-forward ratio of 2:1 to calculate the total FLOPs for active layers. We use the FlopCountAnalysis tool in the fvcore [48] library for calculating the FLOPs of the forward pass. We only consider a single input sample for FLOPs calculation.
A.2 Data Distribution
Fig. 16 illustrates the data distribution among clients for different $\beta$ values used in Section 5.6. Lower $\beta$ values correspond to increasingly heterogeneous data distributions among clients.
A.3 Self-supervised Learning
Algorithm 3 describes the detailed training procedure of MoCo v3 [11].
Algorithm 3 Self-supervised learning (SSL)
Input: encoder $F$, projection head $H$, prediction head $P$, dataset $D$, number of epochs $E$, momentum $\mu$, temperature $\tau$
Output: encoder $F$, projection head $H$, prediction head $P$
1:target branch: $F_{k}\leftarrow F$ and $H_{k}\leftarrow H$
2:for epoch $e=1,2,\dots,E$ do
3:     for each batch $x\in D$ do
4:         $x_{1}\leftarrow\text{Augment}(x)$ and $x_{2}\leftarrow\text{Augment}(x)$
5:         $q_{1}\leftarrow P(H(F(x_{1})))$ and $q_{2}\leftarrow P(H(F(x_{2})))$
6:         $k_{1}\leftarrow H_{k}(F_{k}(x_{1}))$ and $k_{2}\leftarrow H_{k}(F_{k}(x_{2}))$
7:         $\ell_{\text{con}}\leftarrow\ell(q_{1},k_{2},\tau)+\ell(q_{2},k_{1},\tau)$triangleright Eq. (2)
8:         $F,H,P\leftarrow\text{Update}(\nabla\ell_{\text{con}})$
9:         $F_{k},H_{k}\leftarrow\text{Momentum\_Update}(F,H,\mu)$
10:return $F,H,P$
Figure 15: Comparison between FedMoCo, FedMoCo-LW, and Prog-FedSSL at stage $s$.
(a) $\beta=50$
(b) $\beta=5$
(c) $\beta=0.5$
(d) $\beta=0.1$
Figure 16: Client data distribution with different $\beta$ values for the Tiny ImageNet dataset. A darker color denotes a higher number of data samples for a specific class in a client.
Appendix B Additional Experiments
B.1 Weight of Representation Alignment
In LW-FedSSL, the weight term $\alpha$ controls the regularization strength of the representation alignment mechanism in the local training process. We explore different $\alpha$ values in Table 5. Based on the results, we set the default value of $\alpha$ to 0.01.
Table 5: Linear evaluation results of LW-FedSSL with different $\alpha$ values.
$\alpha$ CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
0.01 61.83 32.25 55.15 17.26
0.1 59.21 30.53 52.25 16.14
1 56.92 27.13 52.34 13.00
10 55.84 24.00 50.67 11.66
Table 6: Linear evaluation results with and without weight transfer (WT).
WT CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
FedMoCo-LW w/o 49.38 25.15 46.41 14.35
w/ 50.09 27.16 47.95 14.64
LW-FedSSL w/o 63.27 31.84 53.30 18.37
w/ 61.83 32.25 55.15 17.26
Prog-FedSSL w/o 65.07 39.53 58.79 23.74
w/ 65.64 39.52 60.91 22.43
B.2 Weight Transfer
For FedMoCo-LW, LW-FedSSL, and Prog-FedSSL, we transfer the weights of the prior layer, $L_{s-1}$, to the subsequent layer, $L_{s}$, at the beginning of a new stage $s$. Here, we investigate the impact of weight transfer on the performance of these approaches. Table 6 demonstrates that weight transfer can improve the linear evaluation performance of the baseline FedMoCo-LW across different downstream datasets. For LW-FedSSL, weight transfer improves the performance on STL-10 and CIFAR-100 but decreases on CIFAR-10 and Tiny ImageNet datasets. Similarly, in the case of Prog-FedSSL, weight transfer leads to improved performance on STL-10 but a slight decrease on Tiny ImageNet, while maintaining comparable performance on other datasets. Notably, weight transfer consistently improves performance across all approaches on the STL-10 dataset, a case where the downstream data originates from a similar distribution as the unlabeled pre-training data within clients. This highlights a potential utility of weight transfer in the context of layer-wise training, where downstream data and pre-training data originate from a similar distribution (e.g., semi-supervised scenarios where only a portion of the data is labeled).
B.3 Hidden Layer Dimension in MLP Heads
Following [11], we set the dimension of hidden layers in the projection head $H$ and the prediction head $P$ to 4096 by default. Tables 7 and 8 describe the detailed structures of $H$ and $P$, respectively. We further explore various hidden dimensions, specifically 128, 256, 512, and 1024. The results in Fig. 17 indicate that Prog-FedSSL is relatively sensitive to the hidden layer dimension of the MLP heads.
Table 7: Projection head $H$.
Layer
Input
dimension
Output
dimension
Batch
norm
ReLU
1 192 4096
2 4096 4096
3 4096 256
Table 8: Prediction head $P$.
Layer
Input
dimension
Output
dimension
Batch
norm
ReLU
1 256 4096
2 4096 256
Figure 17: Hidden layer dimension in the MLP heads. Linear evaluation results are shown for the STL-10 dataset.